# Ambiguous Online Learning

### v3

@@ Appendix H Lower Bound for Many Labels
Specifically, the strategy works roughly as follows: on every round, we choose a set $B$ of $l$ hypotheses, out of the set AA of unfalsified hypotheses, by picking the hypotheses $j\in A$ with the least number of prior mistakes $m(j)$. Fixing a bijection $f$ between $B$ and $l+1]\setminus\{0\}$, we let every $j\in B$ predict the label set $l+1]\setminus\{f(j)\}$. Every hypothesis in $A\setminus B$ predicts the full label set $Y=[l+1]$. If the learner predicts $\alpha\subseteq[l+1]$ then a label $y\in[l+1]\setminus\alpha$ is selected, and all surviving hypotheses in AA gain 1 mistake, while at most one hypothesis is falsified. If the learner predicts $\alpha=[l+1]$ then the label $y=0$ is selected, nothing is falsified and all hypotheses in $B$ gain a mistake.
@@
of Proposition 7.5.
Given any finite non-empty set AA and $m:A\to\mathbb{N}$, we recursively construct a tree $T^{(m)}=(V^{(m)},v^{(m)}_{0},P^{(m)})$. It will be equipped with $E_{0}^{(m)}\subseteq E^{(m)}$ and a labeling $\mathbf{y}^{(m)}:E^{(m)}\to[l+1]$ satisfying the usual conditions for ambiguous shattered trees. It will also be equipped with a labeling $\mathbf{h}^{(m)}:V_{f}^{(m)}\to A$.

### v4

@@ Appendix H Lower Bound for Many Labels
Specifically, the strategy works roughly as follows: on every round, we choose a set $B$ of $l$ hypotheses, out of the set $A$ of unfalsified hypotheses, by picking the hypotheses $j\in A$ with the least number of prior mistakes $m(j)$. Fixing a bijection $f$ between $B$ and $l+1]\setminus\{0\}$, we let every $j\in B$ predict the label set $l+1]\setminus\{f(j)\}$. Every hypothesis in $A\setminus B$ predicts the full label set $Y=[l+1]$. If the learner predicts $\alpha\subseteq[l+1]$ then a label $y\in[l+1]\setminus\alpha$ is selected, and all surviving hypotheses in $A$ gain 1 mistake, while at most one hypothesis is falsified. If the learner predicts $\alpha=[l+1]$ then the label $y=0$ is selected, nothing is falsified and all hypotheses in $B$ gain a mistake.
@@
Proof of Proposition 7.5.
Given any finite non-empty set $A$ and $m:A\to\mathbb{N}$, we recursively construct a tree $T^{(m)}=(V^{(m)},v^{(m)}_{0},P^{(m)})$. It will be equipped with $E_{0}^{(m)}\subseteq E^{(m)}$ and a labeling $\mathbf{y}^{(m)}:E^{(m)}\to[l+1]$ satisfying the usual conditions for ambiguous shattered trees. It will also be equipped with a labeling $\mathbf{h}^{(m)}:V_{f}^{(m)}\to A$.